Ben completes a 20km run in 1 hour and 17.0 minutes- an average speed of 4.30m/s. Suppose that during half of the distance Ben ran at 11.0m/s. How fast did he run during the other half?


1
Expert's answer
2021-09-18T14:59:23-0400

The average speed is total distance over total time:


"v_a=\frac DT=\frac{D}{D/(2v)+D/(2u)}=\frac{2vu}{v+u},\\\space\\
u=\frac{vv_a}{2v-v_a}=\frac{11·4.3}{2·11-4.3}=2.67\text{ m/s}."